# Integration of a high-fidelity model of quantum sensors with a map-matching filter for quantum-enhanced navigation

### v1

@@ I Introduction
Navigating accurately and reliably in environments where traditional positioning systems fail is a critical challenge in both civilian and military applications. Enabling resilient, continuous and precise tracking could enhance safety and efficiency, while also facilitating autonomous navigation systems capable of responding in real-time without human intervention. Conventional navigation methods, such as position fixing via the Global Navigation Satellite System (GNSS), are vulnerable to interference, spoofing and jamming and require a line of sight to satellites, making them unsuitable for denied environments. Inertial Navigation Systems (INS), which estimate position, velocity, and orientation by integrating measurements of acceleration and angular velocity from accelerometers and gyroscopes, are widely used in these scenarios due to their ability to operate independently of external signals. However, their performance degrades over time due to the accumulation of measurement and integration errors, resulting in significant drift Jekeli (2005). Mitigating these drifts has led to the development of position-fixing methods, which use external sources to perform periodic corrections and maintain navigational accuracy Gleason (1995); Jekeli (2006). Alternative navigation via gravity gradient map-matching, where observations of the local gravity gradient at different locations are matched with existing database, is particularly promising as an autonomous, stable and passive position fixing method, as gravity is ubiquitous on Earth and immune to jamming or spoofing Welker et al. (2013); Wang et al. (2016a, 2017); Wu et al. (2017); Phillips et al. (2022). Although gravity map-matching has been demonstrated using classical accelerometers Wang et al. (2016b, 2017), it remains challenging due to the small signals of the gravity variations to detect.
@@
Quantum sensors based on cold-atom interferometry Riehle et al. (1991); Kasevich and Chu (1991, 1992), which have evidenced unprecedented performance levels in measuring inertial quantities Gustavson et al. (1997); Canuel et al. (2006); Barrett et al. (2016a); Yankelev et al. (2019); Geiger et al. (2020); Salducci et al. (2024) in laboratories, including gravitational acceleration Peters et al. (1999, 2001); de Angelis et al. (2008); Lellouch et al. (2022), have emerged as promising candidates for alternative navigation. By using atoms, which are immune to manufacturing defects, wear and tear, as test masses, they not only bring improvement in accuracy over classical sensors Peters et al. (1999, 2001); Hu et al. (2013), but also more practicality as they do not require as frequent calibration to retain their long-term stability Durfee et al. (2006); Gillot et al. (2014). Their development has prompted successful demonstrations of quantum gravimetry in the field including on trucks Wang et al. (2022), air Bidel et al. (2020) and marine Bidel et al. (2018) platforms. Additionally, operating them in a gradiometric configuration where a common laser beam is used to extract information from two spatially-separated atom interferometers derives additional benefits through common-mode noise rejection McGuirk et al. (2002); Sorrentino et al. (2014); Biedermann et al. (2015). Common-mode suppression of 140 dB has been demonstrated in laboratories McGuirk et al. (2002) before being recently achieved in practical settings, which, together with 600-fold reduction in tilt susceptibility over a perfect gravimeter, has facilitated successful trials in real-life environments Stray et al. (2022). While this progress has triggered an intense effort to bring quantum gravity gradiometers to practical use in real application contexts Schmidt et al. (2011); Ménoret et al. (2018); Bongs et al. (2019); Wu et al. (2019); Antoni-Micollier et al. (2022); Janvier et al. (2022), such systems are still in an early development stage compared to classical sensors Jekeli (2005); Fang and Qin (2012).
@@
In particular in navigation applications, quantum sensors are required to operate at high performance in entirely different conditions than they are in lab-based environments. Constraints of size, weight and power (SWAP), and bandwidth are often competing with sensitivity requirements. Additionally, platform motion and influences from the external environment generally lead to significant noise and systematic effects Bidel et al. (2020); Barrett et al. (2016a, b); Saywell et al. (2023); de Castanet et al. (2024), whose impact on sensor accuracy, and subsequently on the overall performance of a full gravity-aided navigation system, is yet to be fully quantified.
This lack of understanding is hindering the rapid development of quantum sensors in navigation settings.
In particular, comprehensive sensor data would be key for training and optimizing navigation filters to the specificities of each platform and sensing capability.
With trials being scarce and costly, modelling is an approach of choice.
@@
Significant progress has been made to understand the various contributions to the interferometric phase shift Peters et al. (2001); Barrett et al. (2016b), including understanding errors arising from dynamic conditions de Castanet et al. (2024). However, accurately capturing the full scope of interactions and cross-couplings at the microscopic level remains a significant challenge.
@@
In this paper, we demonstrate the integration of a high-fidelity, microscopic model of a quantum sensor with a map-matching filter, enabling end-to-end simulation of a complete position-fixing solution. Our cold-atom gravity gradiometer model captures a broad spectrum of factors relevant to navigation, accounting for diverse operational environments, including maritime and aerial platforms, as well as the ability to operate with a full range of sensor specifications. Through our simulations, we identify key challenges that must be addressed to advance the use of quantum sensors in navigation. This work constitutes a first significant step towards the future platform-specific optimization of quantum sensors and navigation filters.
@@ II.1 Physics of an atom-interferometry based gravity gradiometer
The measurement process is depicted on Fig. 1. Within each atom interferometric chamber, previously cooled down atoms are addressed by a sequence of three light pulses while they free-fall in the gravity field. A first pulse places the atoms in a quantum superposition of two momentum states, resulting in two wavepackets moving along two separate trajectories. After a time of free-fall, a pulse is applied that redirects the wavepackets towards each other. The atoms are let to free-fall again for a duration before a final pulse is applied that overlaps the trajectories. Within each chamber, the atomic state populations display interference fringes as a function of the accumulated phase difference between the two arms of the interferometer. For atoms in free-fall in the gravity field , this interferometric phase is given by
@@
Additionally, by affecting the motion of the atoms in the beam, inertial effects can cause imperfect overlap of the two wavepackets and impact the fidelity of atom-light pulses, inducing both contrast loss and additional phase shifts [see Appendix B.2].
@@
A further level of cross-couplings arise from the inhomogeneous nature of atomic cloud, comprised of typically atoms, with an approximately Gaussian cloud shape. The individual position of each atom within the beam, which determines the local intensity it experiences, and its individual velocity, which determines the laser frequency it sees as shifted by the Doppler effect, result in each atom experiencing the and pulses differently. Even in static systems, these inter-atom dephasings result in contrast loss and an apparent phase shift at the output of the interferometer Szigeti et al. (2012); Butts et al. (2013); Dunning et al. (2014); Lellouch et al. (2023). In dynamic conditions, inter-atom dephasings introduce cross-coupling effects between the atomic cloud and inertial forces, as each atom inside the cloud experiences not only light fields, but also inertial forces, in a different way. This results in a highly non-trivial interferometric output after cloud-averaging occurs at detection stage. In practice, cross-couplings between the external environment and inhomogeneous clouds can severely degrade the accuracy and sensitivity of measurements, possibly resulting in systematic effects blurring out the signal of interest Geiger et al. (2011); Bidel et al. (2018, 2020); Saywell et al. (2023).
Cross-coupling effects are notoriously challenging to describe and quantify accurately. Detailed phase shift models have been used in the literature, allowing to successfully account for the effects of tilts and rotations de Castanet et al. (2024) and to capture inertial phase shifts via integrating atomic kinematics into sensitivity function analysis Barrett et al. (2016b) 111In Barrett et al. (2016b), contrast loss was further estimated using a simple approximation.. However, phase shift models, which disregard both the effects of pulse infidelities and inter-atom dephasings [see Appendix B.2], do not capture the entirety of cross-couplings effects at play. An exact treatment of these effects requires consideration of the microscopic nature of atomic clouds and of all individual atomic trajectories -which forms the basis of our modelling approach.
@@ II.2 Modelling concept and implementation
To capture the complexity of these mechanisms, we have developed an atom interferometer model that simulates the coupled dynamics (external motion and internal state dynamics) of each individual atom within the atomic cloud when subjected to an arbitrary sequence of light pulses and arbitrary platform dynamics. We adopt a semi-classical approach that treats external motion classically, while internal dynamics are treated quantum mechanically. The state of each atom is parameterised by six external degrees of freedom, namely its three position and three velocity components, and two internal degrees of freedom, namely the coefficients and in the decomposition of its wavefunction onto the two relevant atomic states and , . As the interferometer couples internal and momentum states, these eight degrees of freedom get highly coupled during the interferometric sequence. For instance, the evolution of the atomic state under a light pulse depends on the position and velocity of the atom relative to the beam, while the subsequent evolution of the atom’s position and velocity is influenced by any momentum kick imparted by the light pulse to one part of the atomic wavepacket.
The general principle of our simulation is depicted on Fig. 2. At time , the cloud is assumed to be in a specified state characterized by a position (e.g. Gaussian), velocity (e.g. thermal Maxwell-Boltzmann) and internal state distribution (e.g. all atoms in the ground state). All atoms are treated independently in the simulation. Performing a Monte-Carlo sampling of the initial cloud, we then propagate the external and internal dynamics of each atom along the entire interrogation sequence. This is achieved by splitting the sequence in periods of pulse interaction and periods of free propagation, considering all possible paths in the interferometer.
@@
At each pulse, instantaneous atomic kinematic parameters and at the time t the pulse starts, together with the knowledge of external and laser fields, are used to determine the local light fields that the atom experiences (Rabi frequency , detuning , phase ), taking into account the laser intensity and phase profiles and the Doppler effect. Those fields are used to numerically propagate the equations of motion describing the interaction between the light and the two-level atom. In this work, we have considered a two-photon Raman interferometer, for which the equations of motion take the form of the Schrodinger equation,
@@
Here, denotes the local gravity acceleration while and respectively denote the time-dependent acceleration and rotation vector of the interferometer in the Earth’s frame. The third, fourth and fifth terms in Eq. 2 respectively represent the Coriolis, centripetal, and Euler’s acceleration. As no light coupling is applied between pulses, the time evolution of internal degrees of freedom amounts to a trivial phase.
This process is handled for all possible paths in the interferometer and the final atomic populations in each quantum state are inferred. These populations are then averaged over the initial cloud distribution. As the simulation treats all atoms independently, numerical parallelization is employed. Interferometric fringes can then be obtained by repeating the entire simulation while chirping the laser frequency. By considering two vertically separated interferometers, we have straightforwardly extended this model into a gravity gravimeter model. Typical model outputs include interferometric fringes and ellipses as displayed on Fig. 1, from which gravity and/or gravity gradients can be extracted. This digital model is capable of providing ab initio simulations of gravity gradient measurements from the knowledge of user-specified sensor specifications (including cloud and beams characteristics, sequence details) and conditions of operation (including arbitrary platform dynamics), with no free parameter. Since the simulation accurately replicates the microscopic dynamics taking place in reality, its bottom-up nature inherently allows it to capture cross-coupling effects as a consequence of their microscopic origins.
@@ III.1 Navigation scenarios
Trajectory 1, depicted on Fig. 3(b), describes the typical motion of a ship navigating over 5 hours from sea state 4 to sea state 1. Sea state is 4 for , 3 for , 2 for (with a few occurrences of a sea state 3 between and ) and 1 for . It uses real attitude data (yaw, pitch and roll angles), recorded at 100Hz, to capture the effects of waves, including tilts, rotations and rotation change rates. For simplicity, we have assumed the ship moves at a constant speed of 10m/s and neglected its altitude variations. Centripetal and Coriolis acceleration arising from the ship’s frame being non-inertial are fully accounted for in the simulation.
@@
We assume that the position of the sensor on the platform is such that the initial position of the cloud in the bottom interferometer coincides with the centre of rotation of the platform.
@@
The gravity environment is encoded in a gravity gradient map, obtained from real data, which has 930 points in latitude and 1576 in longitude, with a grid spacing of degrees in both directions. Its values (colorbar on Fig. 3) correspond to the local component of the gravity gradient tensor at altitude , after subtraction of its average value on Earth; this so-called gravity gradient anomaly is typically in the range (). We neglect all other components of the gravity gradient tensor when computing the gravity field at an arbitrary position (latitude, longitude, altitude).

### v2

@@ I Introduction
Navigating accurately and reliably in environments where traditional positioning systems fail is a critical challenge in both civilian and military applications. Enabling resilient, continuous and precise tracking could enhance safety and efficiency, while also facilitating autonomous navigation systems capable of responding in real-time without human intervention. Conventional navigation methods, such as position fixing via the Global Navigation Satellite System (GNSS), are vulnerable to interference, spoofing and jamming and require a line of sight to satellites, making them unsuitable for denied environments. Inertial Navigation Systems (INS), which estimate position, velocity, and orientation by integrating measurements of acceleration and angular velocity from accelerometers and gyroscopes, are widely used in these scenarios due to their ability to operate independently of external signals. However, their performance degrades over time due to the accumulation of measurement and integration errors, resulting in significant drift Jekeli (2005). Mitigating these drifts has led to the development of position-fixing methods, which use external sources to perform periodic corrections and maintain navigational accuracy Gleason (1995); Jekeli (2006). Alternative navigation via gravity gradient map-matching, where observations of the local gravity gradient at different locations are matched with existing database, is particularly promising as an autonomous, stable and passive position fixing method, as gravity is ubiquitous on Earth and immune to jamming or spoofing Wang et al. (2016a, 2017); Wu et al. (2017); Phillips et al. (2022). Although gravity map-matching has been demonstrated using classical accelerometers Wang et al. (2016b, 2017), it remains challenging due to the small signals of the gravity variations to detect.
@@
Quantum sensors based on cold-atom interferometry Riehle et al. (1991); Kasevich and Chu (1991, 1992), which have evidenced unprecedented performance levels in measuring inertial quantities Gustavson et al. (1997); Canuel et al. (2006); Barrett et al. (2016a); Yankelev et al. (2019); Geiger et al. (2020); Salducci et al. (2024) in laboratories, including gravitational acceleration Peters et al. (1999, 2001); de Angelis et al. (2008); Lellouch et al. (2022); Tino (2021), have emerged as promising candidates for alternative navigation. By using atoms, which are immune to manufacturing defects, wear and tear, as test masses, they not only bring improvement in accuracy over classical sensors Peters et al. (1999, 2001); Hu et al. (2013), but also more practicality as they do not require as frequent calibration to retain their long-term stability Durfee et al. (2006); Gillot et al. (2014). Their development has prompted successful demonstrations of quantum gravimetry in the field including on trucks Wang et al. (2022), air Bidel et al. (2020) and marine Bidel et al. (2018) platforms. Additionally, operating them in a gradiometric configuration where a common laser beam is used to extract information from two spatially-separated atom interferometers derives additional benefits through common-mode noise rejection McGuirk et al. (2002); Sorrentino et al. (2014); Biedermann et al. (2015). Common-mode suppression of 140 dB has been demonstrated in laboratories McGuirk et al. (2002) before being recently achieved in practical settings, which, together with 600-fold reduction in tilt susceptibility over a perfect gravimeter, has facilitated successful trials in real-life environments Stray et al. (2022). While this progress has triggered an intense effort to bring quantum gravity gradiometers to practical use in real application contexts Schmidt et al. (2011); Ménoret et al. (2018); Bongs et al. (2019); Wu et al. (2019); Antoni-Micollier et al. (2022); Janvier et al. (2022), such systems are still in an early development stage compared to classical sensors Jekeli (2005); Fang and Qin (2012).
@@
In particular in navigation applications, quantum sensors are required to operate at high performance in entirely different conditions than they are in lab-based environments. Constraints of size, weight and power (SWAP), and bandwidth are often competing with sensitivity requirements. Additionally, platform motion and influences from the external environment generally lead to significant noise and systematic effects Bidel et al. (2020); Barrett et al. (2016a, b); Saywell et al. (2023); d’Armagnac de Castanet et al. (2024), whose impact on sensor accuracy, and subsequently on the overall performance of a full gravity-aided navigation system, is yet to be fully quantified.
This lack of understanding is hindering the rapid deployment of quantum sensors in navigation settings and constitutes a major obstacle to their integration with existing INS. The challenges of integrating quantum sensors into INS architectures are ubiquitous to a range of quantum-enhanced navigation solutions, from quantum gravity-enhanced navigation such as based on map-matching Wang et al. (2016a, 2017); Wu et al. (2017); Phillips et al. (2022) or gravity compensation Welker et al. (2013), to the fusion of classical and quantum inertial capabilities within hybrid inertial sensors Wright et al. (2022); Tennstedt et al. (2023); Wang et al. (2023). In all these examples, effective integration critically depends on the ability to achieve real-time data fusion, including handling calibration systematics, noise and errors on quantum measurements Wright et al. (2022). As these errors become particularly pronounced on moving platforms, how they propagate through INS architectures can be non-trivial. As a result, access to comprehensive sensor datasets is essential for training and optimizing navigation filters to the specificities of each platform and sensor capabilities.
With trials being scarce and costly, modelling is an approach of choice. While classical sensors benefit from decades of modelling and compensation strategies, equivalent models for quantum sensors are still in early development.
@@
Significant progress has been made to understand the various contributions to the interferometric phase shift Peters et al. (2001); Barrett et al. (2016b), including understanding errors arising from dynamic conditions d’Armagnac de Castanet et al. (2024). However, accurately capturing the full scope of interactions and cross-couplings at the microscopic level remains a significant challenge.
@@
In this paper, we demonstrate the integration of a high-fidelity, microscopic model of a quantum sensor with a map-matching filter, enabling end-to-end simulation of a complete 2D position-fixing solution. Our cold-atom gravity gradiometer model captures a broad spectrum of factors relevant to navigation, accounting for diverse operational environments, including maritime and aerial platforms, as well as the ability to operate with a full range of sensor specifications. Through our simulations, we identify key challenges that must be addressed to advance the use of quantum sensors in navigation. This work constitutes a first significant step towards the future platform-specific optimization of quantum sensors and navigation filters.
@@ II.1 Physics of an atom-interferometry based gravity gradiometer
The measurement process is depicted on Fig. 1. Within each atom interferometer, previously cooled down atoms are addressed by a sequence of three light pulses while they free-fall in the gravity field. A first pulse places the atoms in a quantum superposition of two momentum states, resulting in two wavepackets moving along two separate trajectories. After a time of free-fall, a pulse is applied that redirects the wavepackets towards each other. The atoms are let to free-fall again for a duration before a final pulse is applied that overlaps the trajectories. Within each interferometer, the atomic state populations display interference fringes as a function of the accumulated phase difference between the two arms of the interferometer. For atoms in free-fall in the gravity field , this interferometric phase is given by
@@
More generally, cross-couplings between inertial effects and interferometric pulses give rise to complex mechanisms. By affecting the motion of the atoms in the beam, inertial effects directly impact the fidelity of atom-light pulses, inducing both contrast loss and additional phase shifts [see Appendix B.2]. Reciprocally, momentum kicks imparted at each pulse on a given interferometric arm couple with inertial forces to induce complex non-inertial trajectories of the two wavepackets, resulting in their imperfect overlap at the final pulse, further reducing the contrast.
@@
A further level of cross-couplings arise from the inhomogeneous nature of atomic cloud, comprised of typically atoms, with an approximately Gaussian cloud shape. The individual position of each atom within the beam, which determines the local intensity it experiences, and its individual velocity, which determines the laser frequency it sees as shifted by the Doppler effect, result in each atom experiencing the and pulses differently. Even in static systems, these inter-atom dephasings result in contrast loss and an apparent phase shift at the output of the interferometer Szigeti et al. (2012); Butts et al. (2013); Dunning et al. (2014); Lellouch et al. (2023). In dynamic conditions, inter-atom dephasings introduce cross-coupling effects between the atomic cloud and inertial forces, as each atom inside the cloud experiences not only light fields, but also inertial forces, in a different way. This results in a highly non-trivial interferometric output after cloud-averaging occurs at the detection stage. In practice, cross-couplings between the external environment and inhomogeneous clouds can severely degrade the accuracy and sensitivity of measurements, possibly resulting in systematic effects blurring out the signal of interest Geiger et al. (2011); Bidel et al. (2018, 2020); Saywell et al. (2023).
Cross-coupling effects are notoriously challenging to describe and quantify accurately, due to the complex interplay between internal state evolution, external atomic trajectories and matter-light pulse dynamics.
Traditional atom interferometry models rely on the classical action or phase shift formalism to compute and analyse phase differences between interferometer arms using path integral approaches Kasevich and Chu (1992); Storey and Cohen-Tannoudji (1994) or the sensitivity function formalism Gouët et al. (2008).
Such models have been applied in dynamic situations, allowing to successfully account for the effects of tilts and rotations d’Armagnac de Castanet et al. (2024) and to capture inertial phase shifts via integrating atomic kinematics into sensitivity function analysis Barrett et al. (2016b).
However, these approaches typically disregard the effects of pulse infidelities [see Appendix B.2] and neglect recoil-induced trajectory changes. Additionally, they remain based on a single-atom description that cannot capture inter-atom dephasings.
Similarly, contrast loss in dynamic environments has been estimated from modelling the imperfect overlap of atomic wavepackets Barrett et al. (2016b); Roura et al. (2014), but these approaches do not comprehensively consider the microscopic dynamics within the cloud, and therefore do not accurately capture the effects of pulse infidelities, nor how the diversity of atomic trajectories couple with these.
In turn, Monte Carlo approaches have been developed to describe inhomogeneous atomic clouds and account for pulse infidelities arising from spatially-dependent Rabi frequencies, Doppler detunings and decoherence, while concatenated pulse models have allowed to propagate these effects into interferometer contrast models Szigeti et al. (2012); Saywell et al. (2020); Chiarotti et al. (2022). However, these approaches tend to treat pulses independently and do not capture the microscopic interplay between internal state evolution and
individual non-inertial trajectories.
In this work, we capture the range of these mechanisms by combining a microscopic Monte-Carlo model of the atomic cloud with a semi-classical piecewise numerical evolution of the coupled internal-external dynamics. We note that a similar model structure was proposed in Soh et al. (2020), yet did not consider the full range of inertial effects that we do here.
@@ II.2 Modelling concept and implementation
The atom interferometer model that we have developed simulates the coupled dynamics (external motion and internal state dynamics) of each individual atom within the atomic cloud when subjected to an arbitrary sequence of light pulses and arbitrary platform dynamics. We follow a conventional semi-classical approach that treats external motion classically, while internal dynamics are treated quantum mechanically Kasevich and Chu (1992). The state of each atom is parameterised by six external degrees of freedom, namely its three position and three velocity components, and two internal degrees of freedom, namely the coefficients and in the decomposition of its wavefunction onto the two relevant atomic states and , . In the context of the 87Rb two-photon Raman interferometer considered in this work, these states correspond to two hyperfine ground-states, and , respectively associated with different external momenta and with the momentum of the atom. As the interferometer couples internal and momentum states, these eight degrees of freedom get highly coupled during the interferometric sequence. For instance, the evolution of the internal state under a light pulse depends on the position and velocity of the atom relative to the beam, while the subsequent evolution of the atom’s position and velocity is influenced by any momentum kick imparted by the light pulse onto one part of the atomic wavepacket.
The general principle of our simulation is depicted on Fig. 2. At time , the cloud is assumed to be in a specified state characterized by a position (e.g. Gaussian), velocity (e.g. thermal Maxwell-Boltzmann) and internal state distribution (e.g. all atoms in the ground state). All atoms are treated independently in the simulation. Performing a Monte-Carlo sampling of the initial cloud, we propagate the external and internal dynamics of each atom along the entire interrogation sequence. As each atom gets split into wavepackets that experience different non-inertial trajectories, this is achieved, for every atom, by considering all four paths in the interferometer that involve states and (as depicted on the top-right inset of Fig. 1), and splitting the sequence, along each of them, in periods of pulse interaction and periods of free propagation.
@@
At each pulse, instantaneous atomic kinematic parameters and at the time t the pulse starts, together with the knowledge of external and laser fields, are used to determine the local light fields that the atom experiences (Rabi frequency , detuning , phase ), taking into account the laser intensity and phase profiles and the Doppler effect. Those fields are used to numerically propagate the equations of motion describing the interaction between the light and the two-level atom. In this work, we have considered a 87Rb two-photon Raman interferometer, for which the equations of motion take the form of the Schrodinger equation,
@@
Here, denotes the gravity acceleration instantaneously experienced by the atom given its position, while and respectively denote the time-dependent acceleration and rotation vector of the interferometer in the Earth’s frame. The first term, , is calculated from the knowledge of the instantaneous position of the atom , the local gravity on the ground and the gravity gradient tensor [see details in Sec. III.1]; that way, gravity gradient effects within the interferometer are fully included. The third, fourth and fifth terms in Eq. 2 respectively represent the Coriolis, centripetal, and Euler’s acceleration. As no light coupling is applied between pulses, the time evolution of internal degrees of freedom amounts to a trivial phase.
This simulation process is handled for all four paths of the interferometer [see Fig. 2], allowing to track the individual atomic positions, velocities and internal states along each of them. By accurately replicating the microscopic dynamics taking place in reality, it inherently captures cross-coupling effects as a consequence of their microscopic origins, fully accounting for effects like imperfect overlap of the atomic wavepackets and pulse infidelities. After the final pulse, atomic populations in each quantum state are inferred and averaged over the initial cloud distribution. As the simulation treats all atoms independently, numerical parallelization is employed. Interferometric fringes can then be obtained by repeating the entire simulation for different values of the chirp rate (see Appendix A) such that the interferometer phase is scanned. By considering two vertically separated interferometers, we have straightforwardly extended this model into a gravity gradiometer model. Typical model outputs, as displayed on Fig. 1, include interferometric fringes, from which gravity can be extracted, and ellipses, whose fitting allows extraction of gravity gradients. This digital model is capable of providing ab initio simulations of gravity gradient measurements from the knowledge of user-specified sensor specifications (including cloud and beams characteristics, sequence details) and conditions of operation (including arbitrary platform dynamics), with no free parameter.
The model has been quantitatively benchmarked against instrument data collected in laboratory conditions [see for instance Fig. 6 in Lellouch et al. (2022)]. It was also used to guide the development of the field gravity gradiometer used in Ref. Stray et al. (2022). A detailed and thorough comparison with field data would however require exact characterisation, including in-situ analysis, of the instrument and external conditions—which is ongoing work.
@@ III.1 Navigation scenarios
Trajectory 1, depicted on Fig. 3(b), describes the typical motion of a ship navigating over 5 hours from sea state 4 to sea state 1, as defined in the World Meteorological Organization (WMO) sea state code. Sea state is 4 for , 3 for , 2 for (with a few occurrences of a sea state 3 between and ) and 1 for . It uses real attitude data (yaw, pitch and roll angles), recorded at 100Hz, to capture the effects of waves, including tilts, rotations and rotation change rates. For simplicity, we have assumed the ship moves at a constant speed of 10m/s and neglected its altitude variations. While for a gravity sensor vertical motion at sea can induce significant phase shifts, it is less relevant to differential gravity gradiometry use cases as it is to high order common to both interferometers, thereby strongly suppressed in the differential measurement (see Appendix B.2). Centripetal and Coriolis acceleration arising from the ship’s frame being non-inertial are fully accounted for in the simulation.
@@
Figure 4 further shows, for each trajectory, the rotation rates and rotational accelerations of the platform along the three platform body axes.
We assume the quantum sensor to be strapdown with the platform, and its position on the platform to be such that the initial position of the cloud in the bottom interferometer coincides with the centre of rotation of the platform.
@@
The gravity environment is encoded in a gravity gradient map, obtained from real data, which has 930 points in latitude and 1576 in longitude, with a grid spacing of degrees in both directions. Its values (colorbar on Fig. 3) correspond to the local component of the gravity gradient tensor after subtraction of its average value on Earth, (); this so-called gravity gradient anomaly is typically in the range . We assume that does not vary with altitude, and we neglect all other components of the gravity gradient tensor when computing the gravity field at an arbitrary position (latitude, longitude, altitude), namely with and the upward vertical unit vector.